1. Given the following table:

US UK

Commodity-X (units/man hour) 24 4

Commodity-Y (units/man hour) 16 8

If U.S exchange 24X for 24Y with U.K.

i. What is the gain to US and UK?

ii. What is the range for mutually beneficial trade?

iii. What is the dollar price of commodity X in the U.S if the wage rate is $48?

iv. What is the pound price of commodity X in the U.K if the wage rate is £8?

v. What is the dollar price of commodity Y in the U.K if £1 = $2?

vi. What is the pound price of commodity Y in the U.S if £1 = $2?



1
Expert's answer
2021-04-06T06:55:27-0400

i. Both US and UK can gain benefits, if US produces commodity X, UK produces commodity Y, and then they trade to each other.

ii. The range for mutually beneficial trade is for US to buy Y for no more than 1.5X, and for UK to buy X for no more than 2Y.

iii. The dollar price of commodity X in the U.S is $48/24 = $2.

iv. The pound price of commodity X in the U.K is £8/4 = £2.

v. The dollar price of commodity Y in the U.K if £1 = $2 is £8/8×2 = $2.

vi. The pound price of commodity Y in the U.S if £1 = $2 is $48/16×0.5 = £1.5.